# On the Convergence of Moral Self-Correction in Large Language Models

### v1

@@ Abstract
Guangliang Liu1thanks: Equal contribution., Haitao Mao211footnotemark: 1thanks: This work does not relate to the author’s position at Amazon., Bochuan Cao3, Zhiyu Xue4, Xitong Zhang1, Rongrong Wang1, Kristen Marie Johnson1 1Michigan State University, 2Amazon 3Pennsylvania State University, 4University of California, Santa Barbara {liuguan5, zhangxit, wangron6, kristenj}@msu.edu maoht@amazon.com, bccao@psu.edu, zhiyuxue@ucsb.com
@@ 1 Introduction
Section 2 presents background for and the logical framework of our mechanistic analysis for the convergence of self-correction. Section 3 shows empirical evidence that the converged performance exists for various tasks. Section 4 and 5 illustrate how the activated latent concept and model uncertainty evolves through self-correction rounds, respectively. Section 6 identifies activated latent concepts, through model uncertainty, as a key factor driving the converged performance of self-correction.
@@ 3 The General Convergence of Intrinsic Self-Correction
Figure 3: The self-correction performance for six different tasks including both language generation tasks and multi-choice tasks. The x-axis represents the self-correction round and the y-axis indicates the performance evaluated on the corresponding task. The performance of self-correction improves as the interaction round progresses and converges eventually. The self-correction performance of the social bias mitigation task and the jailbreak defense task reaches the best performance in the first round and maintains this optimal performance with no modification for the rest of the interaction rounds.
@@
Experimental Settings. The adopted tasks can be categorized into (1) multi-choice QA tasks: social bias mitigation (Parrish et al., 2022), jailbreak defense (Helbling et al., 2023), and visual question answer (VQA) Tong et al. (2024) (2) generation tasks: commonsense generation (Lin et al., 2020), text detoxification (Gehman et al., 2020; Krishna, 2023), and visual grounding Lin et al. (2014). Notably, visual grounding and visual question answer (VQA) are multi-modality tasks requiring an understanding of both vision and language. The considered model in this paper is zephyr-7b-sft-full (Tunstall et al., 2023), a LLM model further fine-tuned on Mistral-7B-v0.1 (Jiang et al., 2023) with instruction-tuning. GPT-4  [^4] https://openai.com/index/gpt-4-research/ is utilized as the backbone vision-language model for vision-language tasks. We consider a multi-round self-correction pipeline in a conversational scenario (as show in Figure 1), and self-correction instructions are utilized per round. The instruction for the first round is concatenated with the original question. The following instructions are appended with the dialogue history as the post-hoc instruction to correct the misbehavior. Following the setting in Huang et al. (2023a), we set the number of self-correction rounds as a constant. We use 10 rounds for text detoxification and commonsense generation, and 5 rounds for other tasks. More experimental details can be found in Appendix C.
@@
In conclusion, the application of multi-round self-correction consistently enhances performance and eventually achieves convergence. These findings suggest that intrinsic self-correction offers convergence guarantees across a variety of tasks. In the following sections, we introduce how the converged performance is related to activated positive concept and reduced model uncertainty.
@@ 4 Latent Concept
In this section, we investigate how the activated latent concept evolves as the self-correction process progresses, building on the approach of identifying latent concepts to understand in-context learning (Xie et al., 2021) and the morality of LLMs (Lee et al., 2024). In this context, a latent concept is regarded as the moral orientation underlying the input. In the context of detoxification, negative or immoral concepts are associated with toxic content, whereas positive or moral concepts correspond to non-toxic outputs. Similarly, in the text detoxification task, concepts include toxicity and non-toxicity. Since this section, we use Mistral-7B in our analysis for two reasons: (1) it has not been exposed to our benchmarks (BBQ and RealToxicity), which some open-source models have seen during instruction tuning; and (2) it demonstrates strong instruction-following capabilities. Mistral-7B is one of the few models that meet both criteria and is widely adopted in prior work.
@@ 5 Model Uncertainty
We adopt the method of semantic uncertainty Kuhn et al. (2022) to estimate uncertainty for language generation tasks, which involves estimating linguistic-invariant likelihoods by the lens of semantic meanings of the text. For multiple-choice QA tasks, we treat LLM predictions as a classification problem and use normalized logits—i.e., the log-likelihoods of each choice (e.g., (a), (b), (c))—as a measure of model uncertainty, following the approach in Guo et al. (2017) and Kadavath et al. (2022). We estimate model uncertainty by self-correction rounds, and pick up four representative social biases from the BBQ benchmark (Parrish et al., 2022).
@@
Figure 5 presents how the model uncertainty changes as the self-correction round progresses. It is worth noting that self-correction performance converges prior to the point at which model uncertainty reaches its minimum (Fig.3 vs. Fig.5), suggesting that even a moderate level of uncertainty can sufficiently reduce linguistic variation in the outputs of LLMs. In Section 6, we will show this phenomenon is driven by the activated concept by self-correction instructions.
Previous studies Yin et al. (2023); Shen et al. (2024) show that large language models generally not calibrated in their generation process. We test the calibration error during the self-correction process inspired by prior studies (Wang et al., 2021; Ao et al., 2023), showing that less uncertainty can reduce calibration errors. We leverage the ECE error Guo et al. (2017) for QA tasks and the Rank-calibration error (RCE) (Huang et al., 2024) for the language generation task.
@@
Figure 6 presents how the calibration error change as the self-correction round progresses. Experimental results indicate that: (1) All the reported tasks demonstrate a trend of converged calibration error as the rounds progress. (2) The ECE error of QA tasks converged at the first or second round, which helps to explain why the self-correction performance of QA tasks (social bias mitigation) converges in the first iteration as shown in Figure 3. (3) The RCE error of generation tasks show convergence since round 6, aligning with the trend of performance curves (text detoxification) reported in Figure 3. The reduced calibration error provides strong evidence for the effectiveness of self-correction.
@@ 6 Dependence Between Latent Concept and Model Uncertainty
7 Related work
Self-correction is the capability of LLMs that allows them to modify their outputs based on instructions or external feedback. Such ability enables LLMs to adjust their responses for improved accuracy, relevance, and coherence, helping LLMs more effective in various applications. Proper-designed self-correction instruction has revealed empirical success in various application scenarios, e.g., machine translation (Chen et al., 2023), code generation (Madaan et al., 2023), social bias mitigation (Schick et al., 2021). Self-correction techniques (Pan et al., 2023) can be roughly categorized into (1) instruction-based, utilizing vanilla natural language instruction and intrinsic self-correction capability of the LLM (2) external-feedback based one, relying on an external verifier to provide external feedback. Our paper focuses on the intrinsic capability of LLM and the instruction-based self-correction techniques while leaving the external ones as important future work. Moreover, our paper shows correlation with Huang et al. (2023a), a recent empirical analysis paper on the self-correction technique. Our paper can provide additional explanation on phenomena found in Huang et al. (2023a), which shows that LLMs struggle to amend their prior responses where the GPT3.5 almost always believes its initial response is correct. We hypothesize such phenomenon is due to the model initial response reach a high certainty with no further modification in the later stage. Huang et al. (2023a) also finds that enhancement attributed to self-correction in certain tasks may stem from an ill-crafted initial instruction that is overshadowed by a carefully-crafted feedback prompt.
Uncertainty estimation is a crucial approach for examining the inner state of machine learning models with respect to an individual sample or a dataset. However, estimating uncertainty of LLMs, in the context of language generation, presents unique challenges due to the exponentially large output space and linguistic variants. To address these challenges, various estimation techniques are proposed, utilizing token-level entropy Huang et al. (2023b), sentence-level semantic equivalence Kuhn et al. (2022), and the distance in the hidden state space Ren et al. (2022). A reliable uncertainty estimation, which provides the belief of LLMs, is identified as a key step towards safe and explainable NLP systems. Notably, our paper does not aim to develop a more faithful and calibrated LLM with unbiased beliefs. Instead, we leverage LLMs’ uncertainty to interpret self-correction. For more discussion on related works, please refer to Appendix A.
@@ 8 Discussions
Liu et al. (2024) empirically demonstrates that intrinsic moral self-correction is superficial, as it does not significantly alter immorality in hidden states. Our study addresses the question of why intrinsic self-correction is still effective despite its superficiality. Given that intrinsic self-correction relies solely on the internal knowledge of LLMs, the conclusion presented in this paper serves as strong evidence supporting the superficial hypothesis. It suggests that, during pre-training, LLMs may have encountered discourses similar to the input (dialogue history + instructions) in the process of self-correction. We exclude reasoning tasks from our analysis due to ongoing debates surrounding the effectiveness of self-correction in reasoning (Huang et al., 2023a). But Xi et al. (2023) demonstrates the converged performance in reasoning tasks. Intrinsic moral self-correction is a practical instance of the Three Laws of Robotics (Asimov, 1942); with this principle we expect LLMs can follow our abstract orders and take harmless actions.
@@ 9 Conclusion & Future Work
Conclusion. In this paper, we validate the convergence phenomenon of intrinsic self-correction across various tasks and LLMs/VLMs, and reveal that the effectiveness of intrinsic self-correction stems from reduced model uncertainty. Specifically, we show empirical evidence and mathematical simulation that the convergence of activated concepts by self-correction instructions drives the model uncertainty towards convergence, therefore motivating LLMs to approach a converged performance.
@@ Appendix A Related work
The instruction-following capability of LLMs is the foundation for self-correction. However, vanilla LLMs may not be good at following instructions from humans Ouyang et al. (2022). To address this issue, recent LLMs have been equipped with instruction tuning techniques Liu et al. (2023); Rafailov et al. (2024); Ouyang et al. (2022), which utilize templates and response pairs in text-to-text format Raffel et al. (2020) and show effectiveness on following instruction to unseen tasks. More recently, advanced instruction tuning techniques Taori et al. (2023); Longpre et al. (2023); Chung et al. (2024) have been developed to acquire labor-free, task-balancing, and large-scale instruction-following data. To quantify the instruction following capability, Hendrycks et al. (2020); Li et al. (2023b) collect datasets towards scalable and cost-effective evaluation methods. To quantify instruction-following capability, datasets for scalable and cost-effective evaluation methods have been conducted Zeng et al. (2023); Wu et al. (2023); Li et al. (2023a), which evaluates on adverserial, counterfactual, and unnatural instruction following scenarios.
Moreover, our paper shows correlation with Huang et al. (2023a), a recent empirical analysis paper on the self-correction technique. Our paper can provide additional explanation on phenomenons found in Huang et al. (2023a). Huang et al. (2023a) finds that LLMs struggle to amend their prior responses where the GPT3.5 0301 version almost always believes its initial response is correct. We hypothesize such phenomenon is due to the model initial response reach a high certainty with no further modification in the later stage. Huang et al. (2023a) also finds that enhancement attributed to self-correction in certain tasks may stem from an ill-crafted initial instruction that is overshadowed by a carefully-crafted feedback prompt.

### v3

@@ Abstract
Guangliang Liu1thanks: Equal contribution.    Haitao Mao211footnotemark: 1thanks: Work was done at Michigan State University.    Bochuan Cao3   Zhiyu Xue4    Xitong Zhang1   Rongrong Wang1   Kristen Marie Johnson1 1Michigan State University    2Amazon    3Pennsylvania State University     4University of California, Santa Barbara {liuguan5, zhangxit, wangron6, kristenj}@msu.edu maoht@amazon.com     bccao@psu.edu     zhiyuxue@ucsb.com
@@ 1 Introduction
Section 2 presents background for and the logical framework of our mechanistic analysis for the convergence of self-correction. Section 3 shows empirical evidence that the converged performance exists for various tasks. Section 4 and 5 illustrate how the activated latent concept and model uncertainty evolve through self-correction rounds, respectively. Section 6 identifies activated latent concepts, through model uncertainty, as a key factor driving the converged performance of self-correction.
@@ 3 The General Convergence of Intrinsic Self-Correction
Figure 3: The self-correction performance for six different tasks, including both language generation tasks and multi-choice tasks. The x-axis represents the self-correction roun, and the y-axis indicates the performance evaluated on the corresponding task. The performance of self-correction improves as the interaction round progresses and converges eventually. The self-correction performance of the social bias mitigation task and the jailbreak defense task reaches the best performance in the first round and maintains this optimal performance with no modification for the rest of the interaction rounds.
@@
Experimental Settings. The adopted tasks can be categorized into (1) multi-choice QA tasks: social bias mitigation (Parrish et al., 2022), jailbreak defense (Helbling et al., 2023), and visual question answer (VQA) Tong et al. (2024) (2) generation tasks: commonsense generation (Lin et al., 2020), text detoxification (Gehman et al., 2020; Krishna, 2023), and visual grounding Lin et al. (2014). Notably, visual grounding and visual question answer (VQA) are multi-modality tasks requiring an understanding of both vision and language. The considered model in this paper is zephyr-7b-sft-full (Tunstall et al., 2023), a LLM model further fine-tuned on Mistral-7B-v0.1 (Jiang et al., 2023) with instruction-tuning. GPT-4  [^4] https://openai.com/index/gpt-4-research/ is utilized as the backbone vision-language model for vision-language tasks.
We consider a multi-round self-correction pipeline in a conversational scenario (as show in Figure 1), and self-correction instructions are utilized per round. The instruction for the first round is concatenated with the original question. The following instructions are appended with the dialogue history as the post-hoc instruction to correct the misbehavior. Following the setting in Huang et al. (2023a), we set the number of self-correction rounds as a constant. We use 10 rounds for text detoxification and commonsense generation, and 5 rounds for other tasks. More experimental details can be found in Appendix C.
@@
In conclusion, the application of multi-round self-correction consistently enhances performance and eventually achieves convergence. These findings suggest that intrinsic self-correction offers convergence guarantees across a variety of tasks. In the following sections, we introduce how the converged performance is related to the activated positive concept and reduced model uncertainty.
@@ 4 Latent Concept
In this section, we investigate how the activated latent concept evolves as the self-correction process progresses, building on the approach of identifying latent concepts to understand in-context learning (Xie et al., 2021) and the morality of LLMs (Lee et al., 2024). In this context, a latent concept is regarded as the moral orientation underlying the input. In the context of detoxification, negative or immoral concepts are associated with toxic content, whereas positive or moral concepts correspond to non-toxic outputs. Similarly, in the text detoxification task, concepts include toxicity and non-toxicity. Since this section, we use Zephyr-7b in our analysis for two reasons: (1) it has not been exposed to our benchmarks (BBQ and RealToxicity), which some open-source models have seen during instruction tuning; and (2) it demonstrates strong instruction-following capabilities. Zephyr-7B is one of the few models that meet both criteria and is widely adopted in prior work.
@@ 5 Model Uncertainty
We adopt the method of semantic uncertainty Kuhn et al. (2022) to estimate uncertainty for language generation tasks, which involves estimating linguistic-invariant likelihoods by the lens of the semantic meanings of the text. For multiple-choice QA tasks, we treat LLM predictions as a classification problem and use normalized logits—i.e., the log-likelihoods of each choice (e.g., (a), (b), (c))—as a measure of model uncertainty, following the approach in Guo et al. (2017) and Kadavath et al. (2022). We estimate model uncertainty by self-correction rounds, and pick up four representative social biases from the BBQ benchmark (Parrish et al., 2022).
@@
Figure 5 presents how the model uncertainty changes as the self-correction round progresses. It is worth noting that self-correction performance converges prior to the point at which model uncertainty reaches its minimum (Fig.3 vs. Fig.5), suggesting that even a moderate level of uncertainty can sufficiently reduce linguistic variation in the outputs of LLMs. In Section 6, we will show that the phenomenon is driven by the activated concept by self-correction instructions.
Previous studies Yin et al. (2023); Shen et al. (2024) show that large language models are generally not calibrated in their generation process. We test the calibration error during the self-correction process inspired by prior studies (Wang et al., 2021; Ao et al., 2023), showing that less uncertainty can reduce calibration errors. We leverage the ECE error Guo et al. (2017) for QA tasks and the Rank-calibration error (RCE) (Huang et al., 2024) for the language generation task.
@@
Figure 6 presents how the calibration error changes as the self-correction round progresses. Experimental results indicate that: (1) All the reported tasks demonstrate a trend of converged calibration error as the rounds progress. (2) The ECE error of QA tasks converged at the first or second round, which helps to explain why the self-correction performance of QA tasks (social bias mitigation) converges in the first iteration, as shown in Figure 3. (3) The RCE error of generation tasks show convergence since round 6, aligning with the trend of performance curves (text detoxification) reported in Figure 3. The reduced calibration error provides strong evidence for the effectiveness of self-correction.
@@ 7 Discussions
Liu et al. (2024) empirically demonstrates that intrinsic moral self-correction is superficial, as it does not significantly alter immorality in hidden states. Our study addresses the question of why intrinsic self-correction is still effective despite its superficiality. Given that intrinsic self-correction relies solely on the internal knowledge of LLMs, the conclusion presented in this paper serves as strong evidence supporting the superficial hypothesis. It suggests that, during pre-training, LLMs may have encountered discourses similar to the input (dialogue history + instructions) in the process of self-correction. We exclude reasoning tasks from our analysis due to ongoing debates surrounding the effectiveness of self-correction in reasoning Huang et al. (2023a). But Xi et al. (2023) demonstrates the converged performance in reasoning tasks. Intrinsic moral self-correction is a practical instance of the Three Laws of Robotics Asimov (1942); with this principle, we expect LLMs can follow our abstract orders and take harmless actions.
@@ 8 Conclusion & Future Work
Conclusion. In this paper, we validate that intrinsic self-correction consistently converges across diverse tasks and different model architectures, including both LLMs and VLMs. We further reveal that its effectiveness is driven by reduced model uncertainty. Specifically, through empirical evidence and task simulations, we show that the convergence of activated concepts induced by self-correction instructions guides model uncertainty toward a stable state, ultimately enabling LLMs to achieve converged performance.